OpenStudy (anonymous):

In a bag there are 4 green, 3 blue and 5 red discs. two discs are drawn at random, and the first disc is not replaced before the second disc is drawn. Find the probability of drawing. (a) a blue disc on the first draw.

OpenStudy (anonymous):

3/(3+4+5)

OpenStudy (anonymous):

isn't there three possibilities where blue can be drawn on the first draw

OpenStudy (anonymous):

if not replaced: 4 green, 3 blue and 5 red discs= total 12 select 2 blue , there are 3 blue (3/12)* (2/11)=6/132 = 1/22

OpenStudy (anonymous):

the answer is 1/4 though

OpenStudy (anonymous):

probability blue on first draw = number of blue discs / total number of discs = 3 /12 = 1/4